# Faster Diffusion via Temporal Attention Decomposition

### v2

@@ 6.2 Improvement over U-Net-based Models
Tgate can also be applied to U-Net-based models [40, 36]. For all settings shown in Table 2, Tgate enhances the performance of the base models in terms of computational efficiency and FID scores. In particular, Tgate works better when the parameter size of the base model increases. In SDXL, Tgate can reduce the latency by half on the commercial GPU card (from 53.187 to 27.412 s). This indicates the effectiveness and scalability of Tgate on U-Net-based diffusion models. Qualitative analysis is given in Fig. 4 and Appendix F, and the evaluation of text alignment is discussed in Appendix H.
@@ 6.3 Improvement over Acceleration Models
Table 3: Computational complexity, latency, and FID using the LCM distilled from SDXL and PixelArt-Alpha.
@@
Table 4: Comparison with Adaptive Guidance [4] on the MS-COCO validation set based on SDXL and Pixart-Alpha.
@@
Improvement over Consistency Model. Tgate is implemented using a distillation-based method [45, 16], namely the latent consistency model (LCM). The LCM distilled from SDXL  [^3] https://huggingface.co/latent-consistency/lcm-sdxl is first used as the base model, and a grid search is performed for different inference steps. Table 4 reveals that the generation performance is improved in fewer inference steps (i.e., four). To incorporate Tgate into the LCM, the cross-attention prediction of the first or second step ($m=1,2$) is cached and reused in the remaining inference steps. Due to limited inference steps, self-attention is not cached. Table 4 shows the experimental results of the LCM models distilled from SDXL and PixArt-Alpha [^4] https://huggingface.co/PixArt-alpha/PixArt-LCM-XL-2-1024-MS. Although the trajectory is deeply compressed into a few steps, Tgate functions well, and further decreases PixArt-based LCM computation. Thus, the MACs and latency are reduced by 10.98% and 6.02%, respectively, with comparable generation results. As Tgate does not incur any training costs, integrating it with consistency models is valuable and promising. As a reasonable blueprint, distillation-based methods require sampling trajectories from the teacher model; contrarily, Tgate can enhance sampling efficiency, thereby accelerating the learning process. This aspect will be further discussed in a future study. The visualization is provided in Appendix F.
Comparison with Adaptive Guidance. Adaptive guidance [4] offers a strategy for the early termination of CFG. The efficiency of Tgate surpasses that of adaptive guidance, as it innovatively caches and reuses attention. On terminating the CFG in PixArt-Alpha, adaptive guidance yields 2.14T MACs/step, whereas Tgate skips cross-attention and further reduces this value to 1.83T MACs/step. This optimization moderately reduces the computational overhead, as shown in Table 4. Notably, recent trends have shifted toward distillation-based techniques [49, 43] to accelerate inference. These methods compress the denoising process into fewer steps, often achieving single-digit iterations. The student model learns to mimic the CFG-based output during distillation; therefore, CFG decreases during inference, rendering adaptive guidance inapplicable. In contrast, Tgate can fill this gap and further accelerate the distillation-based models, as shown in Table 4. Beyond current capabilities, the superior scalability of Tgate is compared with that of adaptive guidance, particularly with increasing input sizes. This scalability feature of Tgate is further explored in Appendix E.
Table 5: Computational complexity, latency, and FID on the MS-COCO validation set using DeepCache.
@@
Table 6: Zero-shot FIDs on the MS-COCO validation set using the base model of SDXL and different noise schedulers [22, 30, 48].
@@
Improvement over DeepCache. Table 6 compares the performance of DeepCache [32] and Tgate based on SDXL. Although DeepCache is more efficient, Tgate outperforms it in terms of generation quality. Tgate is integrated with DeepCache by reusing cross-attention maps, thereby yielding superior results: MACs and latency of 43.868 T and 14.666 s. Remarkably, DeepCache caches the mid-level blocks to decrease computational load, which is specific to the U-Net architecture. However, its generalizability to other architectures, such as the transformer-based architecture, remains underexplored. Beyond DeepCache, Tgate has wider applications and can considerably improve transformer-based diffusion models, as shown in Table 1. Owing to this adaptability, Tgate is a versatile and potent enhancement for various architectural frameworks. Self-attention is fully operational in the experiment conducted herein and aligned with the DeepCache strategy, as it is not included in the first and last blocks.
Improvement over Different Schedulers. The generalizability of Tgate is evaluated on different noise schedulers. As shown in Table 6, three advanced schedulers [22, 30, 48], are considered that could compress the generation process of a diffusion model to 25 inference steps. Results show that Tgate could consistently achieve stable generation performance in all settings, further indicating its potential for a broad application.
@@ 7 Conclusion and Discussion
The cross-attention and self-attention in the inference process of text-conditional diffusion models are empirically analyzed here, offering the following: i) In the first few inference steps, cross-attention is the primary contributor; however, its influence diminishes in later steps. ii) In contrast, self-attention plays a secondary role initially but gains significance as denoising progresses. iii) By caching and reusing attention maps during scheduled inference steps, Tgate reduces computational demands while still achieving competitive outcomes. These findings encourage further analysis of the role of attention in text-conditional diffusion models. The limitations and broader impacts of this study are discussed in Appendices J and K.
@@ Appendix A Additional Experiments for Temporal Analysis of Cross-Attention
Table A1: Zero-shot FIDs on the MS-COCO validation set using the base model SD-2.1 with DPM-Solver [30]. $m$ is the gate step.
@@
Table A2: Zero-shot FIDs on the MS-COCO validation set using the base model SD-2.1 with DPM-Solver [30]. $m$ is the gate step and $n$ is the total inference number.
@@
EulerD [22] 22.507 21.559 47.108
DPMSolver [30] 22.445 21.557 48.376
DDIM [48] 21.235 20.495 53.150
@@ Appendix B Temporal Analysis of Self-Attention
Analytical Tool Unlike cross-attention, the direct removal of self-attention during inference is not a straightforward process, as the performance deteriorates considerably. Thus, determining the specific contributions of each time step is challenging. To address this issue, a novel analytical approach involving caching and reusing features is proposed. As our core premise, if the features of self-attention can be cached and reused across multiple steps without performance decline, they may be considered less critical. Building on this concept, and drawing parallels with cross-attention analysis, two distinct trajectories, $\mathbb{S}^{F}$ and $\mathbb{S}^{L}$, are introduced.
In $\mathbb{S}^{F}$, self-attention features are cached and reused in the fidelity-improving phase using a gate step $m$ and interval $k$. Specifically, after $m$ inference steps, self-attention is reused for $k$ steps, and self-attention prediction is updated once for the next $k$-step reuse cycle. Contrarily, $\mathbb{S}^{L}$ skips self-attention during the initial semantics-planning phase for every $k$ steps after several warm-up steps, typically set at 2. Then, self-attention is fully integrated into inference after $m$ steps. The performance difference between $\mathbb{S}^{F}$ and $\mathbb{S}^{L}$ indicates the contribution of self-attention at different phases.
CNN-based blocks can process spatial features, performing as interference, in the U-Net architecture; therefore, transformer-based architecture, i.e., PixArt, is alternatively used for the analysis herein.
Results and Discussions For a comprehensive analysis, various experiments are conducted using different values of gate step $m$ and interval $k$. As shown in Table A5, the empirical results convincingly show that self-attention plays a vital role in the latter phases of the process. This is evidenced by consistently higher FID score for $\mathbb{S}^{F}$ than that for $\mathbb{S}^{L}$ across all tests.
The observations from the analysis can be summarized as follows:
Appendix C Overall Pipeline
Figure A1: Pipeline of Tgate for accelerating inference. During the semantics-planning phase, cross-attention (CA) is continuously active, whereas self-attention (SA) is applied every $k$ steps following an initial warm-up period to conserve computational resources. In the fidelity-improvement phase, cross-attention is substituted with a caching mechanism, and self-attention remains operational.
Appendix D Implementation Details
Base Models. Several pre-trained models are used in the experiments: Stable Diffusion-1.5 (SD-1.5) [40], SD-2.1 [40], SDXL [36], PixArt-Alpha [7], SVD [2] and OpenSora [23]. Among them, the SD series are based on convolutional neural networks [11, 12, 57, 24, 19, 50, 13] (i.e., , U-Net [41]). Pixart-Alpha and OpenSora work on the transformer [51] (i.e., , DiT[35]). This experimental setting covers several conditional generation tasks, including text-to-image, text-to-video, and image-to-video tasks.
Acceleration Baselines. For a convincing empirical study, Tgate is compared with several acceleration baseline methods: Latent Consistency Model [31], Adaptive Guidance [4], DeepCache [32], and multiple noise schedulers [22, 30, 48]. Tgate is orthogonal to existing methods used to accelerate denoising inference; therefore, it can be trivially integrated to further accelerate this process.
@@ Appendix D Implementation Details
Evaluation Metrics. Similar to a previous study [36], 10k images from the MS-COCO validation set [28] are used to evaluate the zero-shot generation performance. The images are generated using DPM-Solver [30] with a predefined 25 inference steps and resized to 256 $\times$ 256 resolution to calculate the FID [15]. Tgate is also tested on a high-resolution dataset, namely MJHQ [26] to evaluate its aesthetic quality. The generated images are set at a resolution of 1024 $\times$ 1024, with a total of 10k samples. Following the protocol in ELLA [20], we utilize mPLUG-Large [25] to score the generated samples based on predefined questions. For video generation, the prompts from OpenSora-Sample [23] are used. 10 videos per prompt are generated, and their performance is evaluated based on the CLIP score [37]. As the text branch of SVD is unavailable, SDXL is used to create an image from a prompt, which is then input into SVD to produce the corresponding video. To evaluate the efficiency, Calflops [53] is used to count Multiple-Accumulate Operations (MACs) and the number of parameters (Params.). Furthermore, the latency per sample is assessed on a platform equipped with a Nvidia 1080 Ti.
Appendix E Discussion on Scaling Token Length and Resolution
@@
Tgate can improve efficiency by circumventing cross-attention, which motivated us to examine its contribution to the overall computational cost based on the input size. Specifically, Tgate is compared with SD-2.1 w/o CFG per step to determine the computational cost of cross-attention. Note that SD-2.1 w/o CFG is the computational lower bound for existing methods [4], as it can stop CFG early to accelerate diffusion process. As shown in Table A6, MACs are used as a measure of efficiency. In the default SD-2.1 setting, the resolution is set as 768 with a maximum token length of 77. Results show that cross-attention moderately contributes to the total computational load, which increases exponentially with increasing resolution and token lengths. By omitting cross-attention calculations, Tgate considerably mitigates its adverse effects. For example, in an extreme scenario with the current architecture targeting an image size of 2048 and a token length of 4096 $\times$ 77, MACs can be decreased from 13.457 T to 5.191 T, achieving more than two-fold reduction in computation.
One may argue that existing models do not support such high resolutions or token lengths. However, there is an inevitable trend toward larger input sizes [56, 10, 6]. Furthermore, a recent study [27] has shown the difficulty in computing cross-attention on mobile devices, underscoring the practical benefits of our approach.
@@
Appendix F Additional Visualization
Fig. A4 shows the samples generated using different gate steps. Results show that larger gate steps produce generation results more similar to those of the base models without Tgate. Moreover, the generated samples are visualized based on different steps. As shown in Fig. A5, the difference caused by Tgate is invisible. Considering that different configurations have moderate impacts on FIDs, the samples generated by SDXL and PixArt-Alpha are visualized under various settings, as shown in Fig. A2 and Fig. A3. Although some configurations result in increased FIDs, the changes are nearly imperceptible. These results demonstrate the effectiveness of Tgate.
Appendix G Error Bar
Table A7: Generation performance on MJHQ-10k [26] and MS-COCO [28] datasets with different random seeds on PixArt [7] and SDXL [36].

### v3

@@ 6.2 Improvement over U-Net-based Models
Tgate can also be applied to U-Net-based models (Rombach et al., 2022; Podell et al., 2023). For all settings shown in Table 2, Tgate enhances the performance of the base models in terms of computational efficiency and FID scores. In particular, Tgate works better when the parameter size of the base model increases. In SDXL, Tgate can reduce the latency by half on the commercial GPU card (from 53.187 to 27.412 s). This indicates the effectiveness and scalability of Tgate on U-Net-based diffusion models. Qualitative analysis is given in Fig. 6 and Appendix E, and the evaluation of text alignment is discussed in Appendix H.
@@ 6.3 Improvement over Acceleration Models
Table 3: Computational complexity, latency, and FID using the LCM distilled from SDXL Podell et al. (2023) and PixelArt-Alpha Chen et al. (2024b).
@@
Table 4: Comparison with Adaptive Guidance (Castillo et al., 2023) on the MS-COCO validation set based on SDXL and Pixart-Alpha.
@@
Improvement over Consistency Model. Tgate is implemented using a distillation-based method (Schmidhuber, 1992b; Hinton et al., 2015), namely the latent consistency model (LCM). The LCM distilled from SDXL  [^3] https://huggingface.co/latent-consistency/lcm-sdxl is first used as the base model, and a grid search is performed for different inference steps. Table 4 reveals that the generation performance is improved in fewer inference steps (i.e., four). To incorporate Tgate into the LCM, the cross-attention prediction of the first or second step ($m=1,2$) is cached and reused in the remaining inference steps. Due to limited inference steps, self-attention is not cached. Table 4 shows the experimental results of the LCM models distilled from SDXL and PixArt-Alpha [^4] https://huggingface.co/PixArt-alpha/PixArt-LCM-XL-2-1024-MS. Although the trajectory is deeply compressed into a few steps, Tgate functions well, and further decreases PixArt-based LCM computation. Thus, the MACs and latency are reduced by 10.98% and 6.02%, respectively, with comparable generation results. As Tgate does not incur any training costs, integrating it with consistency models is valuable and promising. As a reasonable blueprint, distillation-based methods require sampling trajectories from the teacher model; contrarily, Tgate can enhance sampling efficiency, thereby accelerating the learning process. This aspect will be further discussed in a future study. The visualization is provided in Appendix E.
Comparison with Adaptive Guidance. Adaptive guidance (Castillo et al., 2023) offers a strategy for the early termination of CFG. The efficiency of Tgate surpasses that of adaptive guidance, as it innovatively caches and reuses attention. On terminating the CFG in PixArt-Alpha, adaptive guidance yields 2.14T MACs/step, whereas Tgate skips cross-attention and further reduces this value to 1.83T MACs/step. This optimization moderately reduces the computational overhead, as shown in Table 4. Notably, recent trends have shifted toward distillation-based techniques (Song et al., 2023; Salimans & Ho, 2022) to accelerate inference. These methods compress the denoising process into fewer steps, often achieving single-digit iterations. The student model learns to mimic the CFG-based output during distillation; therefore, CFG decreases during inference, rendering adaptive guidance inapplicable. In contrast, Tgate can fill this gap and further accelerate the distillation-based models, as shown in Table 4. Beyond current capabilities, the superior scalability of Tgate is compared with that of adaptive guidance, particularly with increasing input sizes. This scalability feature of Tgate is further explored in Appendix D.
Table 5: Computational complexity, latency, and FID on the MS-COCO validation set using DeepCache (Ma et al., 2024).
@@
Table 6: Zero-shot FIDs on the MS-COCO validation set using the base model of SDXL and different noise schedulers (Karras et al., 2022; Lu et al., 2022; Song et al., 2021).
@@
Improvement over DeepCache. Table 6 compares the performance of DeepCache (Ma et al., 2024) and Tgate based on SDXL. Although DeepCache is more efficient, Tgate outperforms it in terms of generation quality. Tgate is integrated with DeepCache by reusing cross-attention maps, thereby yielding superior results: MACs and latency of 43.868 T and 14.666 s. Remarkably, DeepCache caches the mid-level blocks to decrease computational load, which is specific to the U-Net architecture. However, its generalizability to other architectures, such as the transformer-based architecture, remains underexplored. Beyond DeepCache, Tgate has wider applications and can considerably improve transformer-based diffusion models, as shown in Table 1. Owing to this adaptability, Tgate is a versatile and potent enhancement for various architectural frameworks. Self-attention is fully operational in the experiment conducted herein and aligned with the DeepCache strategy, as it is not included in the first and last blocks.
Improvement over Different Schedulers. The generalizability of Tgate is evaluated on different noise schedulers. As shown in Table 6, three advanced schedulers (Karras et al., 2022; Lu et al., 2022; Song et al., 2021), are considered that could compress the generation process of a diffusion model to 25 inference steps. Results show that Tgate could consistently achieve stable generation performance in all settings, further indicating its potential for a broad application.
Additional Comparison with other methods. Beyond the previously discussed methods, we explore integrating our approach with other accelerated diffusion models: (1) SSD (Gupta et al., 2024), which reduces model size via distillation from a larger model, and (2) ToMe (Bolya & Hoffman, 2023), which accelerates inference by compressing token counts. Further details are provided in Appendix G.
@@ 7 Conclusion and Discussion
The cross-attention and self-attention in the inference process of text-conditional diffusion models are empirically analyzed here, offering the following: i) In the first few inference steps, cross-attention is the primary contributor; however, its influence diminishes in later steps. ii) In contrast, self-attention plays a secondary role initially but gains significance as denoising progresses. iii) By caching and reusing attention maps during scheduled inference steps, Tgate reduces computational demands while still achieving competitive outcomes. These findings encourage further analysis of the role of attention in text-conditional diffusion models.
Limitations
We acknowledge the challenge of further improving a distilled diffusion model when working with few inference steps. In such cases, Tgate can only offer a 10% reduction in computation cost (MACs) for models with highly compressed inference steps, such as LCM. However, considering that a model may be used billions of times daily, this reduction is significant. The proposed Tgate is particularly valuable since it does not require additional training costs or result in significant performance drops. Furthermore, Tgate can also be integrated into the distillation process, where it accelerates the generation process of the teacher model, potentially speeding up the training of the student model.
Empirical studies suggest that using Tgate may cause a slight decline in text-image alignment performance (Sec. H) but generally improves the FID score (Sec. 6.1). Based on visualizations, we speculate that although TGATE-generated images are similar to those without it, they tend to produce simpler patterns and objects, akin to outcomes seen in other acceleration methods. To address this, we provide a set of parameters ($k$ and $m$) to balance efficiency and performance, allowing adaptation to different application scenarios. Despite these trade-offs, Tgate reduces inference time by nearly half, making it an effective solution for most cases.
Broader Impacts
@@ Acknowledgment
Jinheng Xie and Mike Shou are only supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Award No: MOE-T2EP20124-0012).
@@ Appendix A Additional Experiments for Temporal Analysis of Cross-Attention
Table A1: Zero-shot FIDs on the MS-COCO validation set using the base model SD-2.1 with DPM-Solver (Lu et al., 2022). $m$ is the gate step.
@@
Table A2: Zero-shot FIDs on the MS-COCO validation set using the base model SD-2.1 with DPM-Solver (Lu et al., 2022). $m$ is the gate step and $n$ is the total inference number.
@@
EulerD (Karras et al., 2022) 22.507 21.559 47.108
DPMSolver (Lu et al., 2022) 22.445 21.557 48.376
DDIM (Song et al., 2021) 21.235 20.495 53.150
@@ Appendix B Temporal Analysis of Self-Attention
For a comprehensive study, we track the generation performance using different values of gate step $m$ and interval $k$. As shown in Table A5, the empirical results convincingly show that self-attention plays a vital role in the latter phases of the process. This is evidenced by consistently higher FID score for $\mathbb{S}^{F}$ than that for $\mathbb{S}^{L}$ across all tests.
Appendix C Implementation Details
Base Models. Several pre-trained models are used in the experiments: Stable Diffusion-1.5 (SD-1.5) (Rombach et al., 2022), SD-2.1 (Rombach et al., 2022), SDXL (Podell et al., 2023), PixArt-Alpha (Chen et al., 2024b), SVD (Blattmann et al., 2023) and OpenSora (Lab & etc., 2024). Among them, the SD series are based on convolutional neural networks (Fukushima, 1979; 1980; Zhang et al., 1988; LeCun et al., 1989; Hochreiter, 1991; Srivastava et al., 2015; He et al., 2016) (i.e., U-Net (Ronneberger et al., 2015)). Pixart-Alpha and OpenSora work on the transformer (Vaswani et al., 2017) (i.e., DiT(Peebles & Xie, 2023)). This experimental setting covers several conditional generation tasks, including text-to-image, text-to-video, and image-to-video tasks.
Acceleration Baselines. For a convincing empirical study, Tgate is compared with several acceleration baseline methods: Latent Consistency Model (Luo et al., 2023), Adaptive Guidance (Castillo et al., 2023), DeepCache (Ma et al., 2024), and multiple noise schedulers (Karras et al., 2022; Lu et al., 2022; Song et al., 2021). Tgate is orthogonal to existing methods used to accelerate denoising inference; therefore, it can be trivially integrated to further accelerate this process.
@@ Appendix C Implementation Details
Evaluation Metrics. Similar to a previous study (Podell et al., 2023), 10k images from the MS-COCO validation set (Lin et al., 2014) are used to evaluate the zero-shot generation performance. The images are generated using DPM-Solver (Lu et al., 2022) with a predefined 25 inference steps and resized to 256 $\times$ 256 resolution to calculate the FID (Heusel et al., 2017). Tgate is also tested on a high-resolution dataset, namely MJHQ (Li et al., 2023) to evaluate its aesthetic quality. The generated images are set at a resolution of 1024 $\times$ 1024, with a total of 10k samples. Following the protocol in ELLA (Hu et al., 2024), we utilize mPLUG-Large (Li et al., 2022) to score the generated samples based on predefined questions. For video generation, the prompts from OpenSora-Sample (Lab & etc., 2024) are used. 10 videos per prompt are generated, and their performance is evaluated based on the CLIP score (Radford et al., 2021). As the text branch of SVD is unavailable, SDXL is used to create an image from a prompt, which is then input into SVD to produce the corresponding video. To evaluate the efficiency, Calflops (xiaoju ye, 2023) is used to count Multiple-Accumulate Operations (MACs) and the number of parameters (Params.). Furthermore, the latency per sample is assessed on a platform equipped with a Nvidia 1080 Ti.
Appendix D Discussion on Scaling Token Length and Resolution
@@
Tgate can improve efficiency by circumventing cross-attention, which motivated us to examine its contribution to the overall computational cost based on the input size. Specifically, Tgate is compared with SD-2.1 w/o CFG per step to determine the computational cost of cross-attention. Note that SD-2.1 w/o CFG is the computational lower bound for existing methods (Castillo et al., 2023), as it can stop CFG early to accelerate diffusion process. As shown in Table A6, MACs are used as a measure of efficiency. In the default SD-2.1 setting, the resolution is set as 768 with a maximum token length of 77. Results show that cross-attention moderately contributes to the total computational load, which increases exponentially with increasing resolution and token lengths. By omitting cross-attention calculations, Tgate considerably mitigates its adverse effects. For example, in an extreme scenario with the current architecture targeting an image size of 2048 and a token length of 4096 $\times$ 77, MACs can be decreased from 13.457 T to 5.191 T, achieving more than two-fold reduction in computation.
One may argue that existing models do not support such high resolutions or token lengths. However, there is an inevitable trend toward larger input sizes (Zhang et al., 2024; Esser et al., 2024; Chen et al., 2024a). Furthermore, a recent study (Li et al., 2024) has shown the difficulty in computing cross-attention on mobile devices, underscoring the practical benefits of our approach.
Figure A3: Illustration of cross-attention map differences between consecutive inference steps. We sample cross-attention maps from various blocks in SD-2.1 to monitor their convergence. The consistent convergence across different blocks supports the generality of our observations.
Figure A4: Illustration of cross-attention maps during different inference steps. We use up_block.1.transformer_block.2.attn2 in SD-2.1 as a representative module to visualize its output for a given prompt, "a delicious salad with beef, stock photo". During inference, the feature map converges rapidly and stabilizes, which aligns with our observations.
@@
Appendix E Additional Visualization
To support Sec. 3, we further analyze the distribution of different blocks’ attention modules. As shown in Fig. A3, we sample the attention maps from down_blocks, mid_block, and up_blocks, and track their changes during different steps. The convergence rates are slightly different, but they share the same trend, which further indicates that this interesting phenomenon is common among different blocks. Meanwhile, as shown in Fig. A4, we also visualize the feature map of up_block as the representative one during different inference steps.
We also provide visualizations of Tgate with different models and configurations. Fig. A5 shows the samples generated using different gate steps. Results show that larger gate steps produce generation results more similar to those of the base models without Tgate. Moreover, the generated samples are visualized based on different steps. As shown in Fig. A6, the difference caused by Tgate is invisible. Considering that different configurations have moderate impacts on FIDs, the samples generated by SDXL and PixArt-Alpha are visualized under various settings, as shown in Fig. A1 and Fig. A2. Although some configurations result in increased FIDs, the changes are nearly imperceptible. These results demonstrate the effectiveness of Tgate.
Table A7: Generation performance on MJHQ-10k (Li et al., 2023) and MS-COCO (Lin et al., 2014) datasets with different random seeds on PixArt (Chen et al., 2024b) and SDXL (Podell et al., 2023).